OpenStudy (anonymous):

differentiate: -pir^2/2

OpenStudy (anonymous):



sam (.sam.):

\[\frac{d}{d r}(-\pi r^2/2)?\]

OpenStudy (anonymous):

yes

sam (.sam.):

Well just apply the power rule \[\large \frac{d}{dx}(x^n)= nx^{n-1}\]

sam (.sam.):

You get \[\frac{d}{d r}(\frac{-\pi r^2}{2}) =-\frac{2\pi r}{2}=-\pi r\]

OpenStudy (anonymous):

ok thanks